### v1

@@ 3.2.2 Few-shot learning
Our experimental results demonstrate the performance capabilities of BrainFounder relative to the baseline model, SwinUNETR, under constrained training data conditions. As depicted in Figure 5, BrainFounder consistently matched the performance of SwinUNETR across higher levels of available training data and outperformed SwinUNETR when training data was constrained to 10% and 20%. When trained with the full 100% data, both models achieved nearly equivalent accuracy. However, as the amount of training data decreased, BrainFounder exhibited superior robustness and adaptability. Notably, at the 20% and 10% data availability levels, BrainFounder maintained higher segmentation accuracy.
Fig. 5: Dice coefficients for baseline (SwinUNETR) and our model across different levels of training data availability.
Overall, the BrainFounder (64M) model provides the best balance between complexity and performance, as evidenced by its leading average Dice coefficient. These results demonstrate the potential benefits of pretraining on the data from a large number of health subjects.
@@ 3.2.3 Modality Restriction
Models trained on T1 images alone demonstrated a significant Dice coefficient reduction of 0.168 when compared to supervised learning with all 4 modalities (mean Dice across folds of 0.8971 vs. 0.7208). However, this reduction is less than that from the same-size SwinUNETR model with the same modality restriction (mean Dice 0.678). Findings for each fold are summarized in Table 5.
BrainFounder (T1 only) SwinUNETR (T1 only) SwinUNETR (T1, T1ce, T2, T2-FLAIR) Fold 1 0.718 0.700 0.885 Fold 2 0.721 0.672 0.906 Fold 3 0.707 0.657 0.898 Fold 4 0.731 0.686 0.892 Fold 5 0.725 0.676 0.904 Average 0.721 0.678 0.897
Table 5: Performance comparison of modality restricted models on the BraTS dataset.
@@ 3.3 Evaluation on ATLAS Challenge Dataset
Our model’s performance on the ATLAS v2.0 dataset was compared against the top-performing models listed on the challenge leaderboard. The results, as summarized in Table 6, demonstrate that our model achieved a Dice score of 0.712, a lesion-wise F1-score of 0.711, simple lesion count of 3.421, and volume difference of 8993.85. These scores would place our model within the top 3 models in the training set leaderboard. Worth noting again is that our training protocol did not include 5-fold cross-validation due to the lack of predetermined folds for with ATLAS can be evaluated on in the training set and the unavailability of some models from the leaderboard for independent validation. Instead, our approach focused on maximizing the comparability with the leaderboard conditions by adhering closely to their reported training setups.
Metric BrainFounder SwinUNETR CTRL(*†) HeRN(*‡) POBOTRI(*) Dice (↑) 0.712 0.703 0.663 0.718 0.663 Lesion-wise F1 Score (↑) 0.711 0.703 0.556 0.724 0.559 Simple Lesion Count (↓) 3.421 3.677 4.657 2.750 4.500 Volume Difference (↓) 8993.85 9165.18 8804.91 6162.00 9535.23
Table 6: Performance comparison of segmentation models on the ATLAS v2.0 dataset. All metrics from the challenge (Dice coefficient, Lesion-wise F1 Score, Simple Lesion Count, and Volume Difference) are included for each model for each model. Scores for models marked with an asterisk (*) are sourced directly from the official challenge leaderboard and pertain to their performance on the training set. CTRL (denoted with †) is the official challenge winner, while HeRN (denoted with ‡) leads on the training set.
@@
The findings from our work with BrainFounder, especially the "Small" model comprising 64 million parameters, signify a noteworthy progression in AI-driven neuroimaging technology. The framework’s novel two-stage pretraining strategy—initially utilizing a broad dataset of multi-modal neuroimages from the generally healthy population found in the UK Biobank, followed by training on pathological brain MRIs from the BraTS dataset—has demonstrated substantial efficacy. This approach enables the framework to acquire a detailed comprehension of normal brain anatomy, a vital aspect for the precise identification of anomalies, like those encountered in brain tumor segmentation tasks.
@@
One of the key findings is the superior performance of the BrainFounder-Small (64M) model over the SwinUNETR model and other BrainFounder variants. Based on our limited explored range of parameters, our model performs best with an intermediate number of parameters. This suggests that an optimal balance of model complexity and training data is crucial. It is also indicative of the importance of large-scale datasets in training AI models for medical imaging, as even the one-stage pretraining model showed significant effectiveness. However, there is still a possibility to see higher performance using a higher number of parameters that we did not explore.
Our results from our study limiting training data indicate that BrainFounders training methods potentially offers better generalization from limited data, a crucial factor for practical applications in medical imaging where annotated data can be scarce. These findings suggest that the enhancements integrated into BrainFounder are effective in optimizing performance under varying data constraints, thereby affirming its suitability for real-world deployment in medical imaging contexts.
@@
Further, the comparable performance of the one-stage 62M model with the two-stage approach indicates that extensive pretraining on a large and diverse dataset like UKB might be sufficient for effective model training, reducing the need for additional pretraining on targeted datasets. This insight could streamline future AI model development in medical imaging, especially in scenarios where specific pathological datasets are limited or hard to acquire.
@@
In our modality restriction experiment, our model sees significant reduction in quality when training with multiple. This difference indicates that the multiple modalities present in BraTS contain important information not present in just T1 images about tumor segmentation. However, BrainFounder’s better performance under these difficult conditions when compared to the base SwinUNETR model validate the feasibility of our extensible approach to handling varying numbers of modalities. Our results demonstrating that the pretraining steps still have a positive effect when the information provided by the original pretraining steps requires truncation. Moreover, our results on ATLAS (discussed more in depth below) further support our method of handling multiple modalities.
@@
The results of our study on the ATLAS dataset indicate BrainFounder’s training scheme is generalizable and effective at more than just tumor segmentation, a trait desirable for foundation models. While methods specifically adapted to optimizing results on this dataset do outperform ours, we still maintain third place in the leaderboard. Remarkably, our results were achieved without the use of ensemble learning techniques, which are commonly employed to boost performance by leveraging the strengths of multiple models. The fact that our single-model approach is competitive with ensemble models underscores the robustness and efficiency of our model in managing the intricacies of medical image analysis. We believe that the methodology used for BrainFounder can be refined and extended to move towards a Medical Foundation Model for neuroimages.
@@
However, it’s important to note that while BrainFounder shows promise in brain tumor segmentation and brain region segmentation, its application in other neuroimaging tasks remains to be explored. One such task is brain tissue segmentation - a common task in automated analysis. Future research should investigate its adaptability to other neurological conditions, its performance in different clinical environments, and its usefulness in additional common analysis tasks.
@@
In conclusion, BrainFounder is a significant step forward in medical AI, particularly neuroimaging. Its development underscores the potential of AI in enhancing diagnostic accuracy and efficiency, paving the way for more advanced, adaptable, and robust AI tools in healthcare.
@@
This work was partially supported by the National Science Foundation (1908299, 2318984). We greatly acknowledge the support from NVIDIA AI Technology (NVAITC) to this research project, especially the great support with GPU parallelization technology. We extend our appreciation to the contributors of the UK Biobank and BraTS datasets, whose extensive data collection efforts have been fundamental to the success of this research. Special thanks to the team at the University of Florida’s HiPerGator-AI supercomputer for providing the computational resources that were essential in training the BrainFounder models.
@@ 6.1 UK Biobank Data
Table 8 provides fold-wise comparison of our BrainFounder models across all tested parameters.
SwinUNETR (62M) BF-T (62M) BF-S (64M) B.F (69M) One-Stage (62M) One-Stage (64M) One-Stage (69M) Fold 1 0.8854 0.9027 0.9032 0.9014 0.9019 0.8994 0.8999 Fold 2 0.9059 0.9181 0.9182 0.9164 0.9188 0.9186 0.9055 Fold 3 0.8981 0.9102 0.9121 0.9097 0.9119 0.9125 0.9002 Fold 4 0.8924 0.9103 0.9100 0.9070 0.9107 0.9133 0.9109 Fold 5 0.9035 0.9139 0.9141 0.9101 0.9132 0.9114 0.9103 Average 0.8971 0.9110 0.9115 0.9089 0.9112 0.9110 0.9054
Table 8: Comparison of BrainFounder models through 5-fold cross-validation with metric Dice coefficient on BraTS. SwinUNETR is the winning solution on BraTS challenge 2021, which is performed with fully supervised learning without UKB pretraining. BrainFounder is the proposed method, which is conducted with the two-stage pretraining and then finetuning on the target dataset. The one-stage means that pretraining on UKB is performed but not on the BraTS. Note: The performance results for SwinUNETR were published on the official GitHub, utilizing hyper-parameter settings similar to those in our finetuning stage but without implementing the ensembling approach that was described in the published work.

### v3

@@ 2.2.2 Few-shot learning
Our experimental results demonstrate the performance capabilities of BrainSegFounder relative to the baseline model, SwinUNETR, under constrained training data conditions. As depicted in Figure 5, BrainSegFounder consistently matched the performance of SwinUNETR across higher levels of available training data and outperformed SwinUNETR when training data was constrained to lower input percentages. As the percentages of training data approached 40% of the input data, both models achieved nearly equivalent accuracy. However, as the amount of training data decreased, BrainSegFounder exhibited superior robustness and adaptability. Notably, at all data availability levels, BrainSegFounder maintained higher mean segmentation accuracy. The results presented in Figure 5 for these input levels are an average of 5 independent subsets of input data to account for variability in small datasets. Supplemental Table 9 contains the results for each of these subsets.
Fig. 5: Dice coefficients for baseline (SwinUNETR) and our model across different levels of training data availability. All models were trained 5 times to account for variability in the input data randomly selected. Error bars represent $\pm$ one standard deviation.
Overall, the BrainSegFounder (64M) model provides the best balance between complexity and performance, as evidenced by its leading average Dice coefficient. These results demonstrate the potential benefits of pretraining on the data from a large number of health subjects.
@@ 2.2.3 Modality Restriction
Our model can effectively adapt with fewer or single modality data in the downstream task. Table 5 shows the five-fold cross-validation and average Dice scores comparing SSL-pretrained BrainSegFounder and fully supervised SwinUNETR, both using T1-weighted MRI only for all training and inference stages. Our multi-stage pretraining demonstrated a DICE score improvement of 0.04 (6%), indicating the substantial benefit of multi-stage pretraining when the downstream task has limited data modality.
BrainSegFounder (T1w only) SwinUNETR (T1w only)
Fold 1 0.718 0.700
Fold 2 0.721 0.672
Fold 3 0.707 0.657
Fold 4 0.731 0.686
Fold 5 0.725 0.676
Average 0.721 0.678
Table 5: Performance comparison of modality restricted models on the BraTS dataset. SwinUNETR is fully supervised learning on T1-weighted MRI without pretraining, while BrainSegFounder uses our multi-stage pretraining on UKB and BraTS T1-weighted MRI and is then finetuned on BraTS T1-weighted MRI.
@@ 2.3 Evaluation on ATLAS Challenge Dataset
Our model’s performance on the ATLAS v2.0 dataset was compared against the top-performing models listed on the challenge leaderboard. The results, as summarized in Table 6, demonstrate that our model achieved a Dice score of 0.712, a lesion-wise F1-score of 0.711, simple lesion count of 3.421, and volume difference of 8993.85. These scores would place our model within the top 3 models in the training set leaderboard. Worth noting again is that our training protocol did not include 5-fold cross-validation due to the lack of predetermined folds for which ATLAS can be evaluated in the training set. Top-performing models on the leaderboard [^1] https://atlas.grand-challenge.org/evaluation/lesion-segmentation-hidden-test-set/leaderboard/ (CTRL, HeRN, and POBOTRI) were not available for independent validation. Instead, our approach focused on maximizing the comparability with the leaderboard conditions by adhering closely to their reported training setups.
Metric BrainSegFounder SwinUNETR CTRL(*†) HeRN(*‡) POBOTRI(*)
Dice (↑) 0.712 0.703 0.663 0.718 0.663
Lesion-wise F1 Score (↑) 0.711 0.703 0.556 0.724 0.559
Simple Lesion Count (↓) 3.421 3.677 4.657 2.750 4.500
Volume Difference (↓) 8993.85 9165.18 8804.91 6162.00 9535.23
Table 6: Performance comparison of segmentation models on the ATLAS v2.0 dataset. All metrics from the challenge (Dice coefficient, Lesion-wise F1 Score, Simple Lesion Count, and Volume Difference) are included for each model for each model. Scores for models marked with an asterisk (*) are sourced directly from the official challenge leaderboard and pertain to their performance on the training set. All of these models utilize ensemble learning methods. CTRL (denoted with †) is the official challenge winner, while HeRN (denoted with ‡) leads on the training set. Top-performing models on the leaderboard (CTRL, HeRN, and POBOTRI) were unavailable for independent validation.
@@
The findings from our work with BrainSegFounder, especially the "Small" model comprising 64 million parameters, signify a noteworthy progression in 3D foundation models for neuroimage segmentation. The framework’s novel two-stage pretraining strategy—initially utilizing a broad dataset of multi-modal neuroimages from the generally healthy population found in the UK Biobank, followed by training on diseased brain MRI volumes from the BraTS dataset—has demonstrated substantial efficacy. The first stage of pretraining enables the framework to capture the latent representation of normal brain anatomy, a vital aspect for the precise identification of anomalies, like those encountered in brain tumor segmentation tasks. The second stage of pretraining learns the spatial distribution and texture representations of lesions presented in different brain disorders.
Table 4 demonstrates superior performance of BrainSegFounder compared to state-of-the-art approaches including SwinUNETR, nn-UNET, SegResNet, TransBTS, and the brain segmentation foundation model published in Model-Zoo. Each of these models represents a significant advancement in brain segmentation techniques. SwinUNETR is the most direct comparison - its architecture is identical to ours, and outperforming SwinUNETR on these tasks indicates that our multi-stage, self-supervised pretraining method is effective in improving segmentation performance. nn-UNeT and SegResNet both were former BraTS challenge winners. As such, they are empirically validated models that excel in the BraTS challenge. By outperforming these models, BrainSegFounder demonstrates its capability to perform segmentation tasks with exceptionally high accuracy and precision. TransBTS utilizes both convolutional and transformer-based architectures to Swinoth local and global context; our superior results compared to TransBTS demonstrate that our model’s performance increase is not merely due to the inclusion of a transformer based architecture and further validate our pipeline. The brain segmentation foundation model from Model-Zoo serves as a comprehensive pre-trained model designed specifically for brain imaging tasks. By demonstrating increased results compared to this novel foundation model, we show that our methodology could streamline the creation of more effective medical foundation models for segmentation.
@@
One of the key findings is the superior performance of the BrainSegFounder-Small (64M) model over other BrainSegFounder variants. Based on our limited explored range of parameters, our model performs best with an intermediate number of parameters. This suggests that an optimal balance of model complexity and training data is crucial. It is also indicative of the importance of large-scale datasets in training 3D vision foundation models for medical imaging, as even the one-stage pretraining model showed significant effectiveness. However, there is still a possibility to see higher performance using a higher number of parameters and large-scale diverse training data that we did not explore in this work.
@@
Further, the comparable performance of the one-stage 62M model with the two-stage approach indicates that extensive pretraining on a large and diverse dataset like UKB might be sufficient for effective model training, reducing the need for additional pretraining on targeted datasets. This insight could streamline future 3D foundation model development for medical imaging, especially in scenarios where specific pathological datasets are limited or hard to acquire.
Our results from our study limiting training data in few-shot learning indicate that BrainSegFounder’s training methods potentially offer better generalization from limited data, a crucial factor for practical applications in medical imaging where annotated data can be scarce. Though only a slight improvement, the BrainSegFounder consistently outperforms the baseline model at lower levels of input data (see Fig 5 and Supplemental Table 9). Even with incredibly limited data, our Stage 2 self-supervised pretraining serves as a meaningful inclusion in the training pipeline. These findings suggest that the enhancements integrated into BrainSegFounder are effective in optimizing performance under varying data constraints, thereby affirming its suitability for real-world deployment in medical imaging contexts.
@@
In our modality restriction experiment, our model sees a significant reduction in quality when training with fewer modalities. This drop indicates that the multiple modalities present in BraTS contain important information not present in just T1-weighted MRI images about tumor segmentation. However, BrainSegFounder’s better performance under these more challenging scenarios with limited modality input when compared to the base SwinUNETR model validates the feasibility of our extensible approach to handling varying numbers of modalities. When fewer modalities are present, our training scheme still provides valuable information and performance improvements by keeping only the layers trained on the modality present. Similarly, our results using all four modalities present in BraTS suggest our method effectively utilizes information given in the pretraining step when presented with additional modalities. Therefore, we conclude that the pretraining steps have a positive effect even when the model is provided with more or less information than is present in the original pretraining stage. Moreover, our results on ATLAS (discussed below) further support our method of handling multiple modalities.
@@
The performance of BrainSegFounder on the ATLAS dataset indicates that its training scheme is generalizable and effective at more than just tumor segmentation, a trait desirable for foundation models. While methods specifically adapted to optimizing results on this dataset do outperform ours, we still maintain third place in the leaderboard. Remarkably, our results were achieved without the use of ensemble learning techniques, which are commonly employed to boost performance by leveraging the strengths of multiple models. The fact that our single-model approach is competitive with ensemble models underscores the robustness and efficiency of our model in managing the intricacies of medical image analysis. We believe that the methodology used for BrainSegFounder can be refined and extended to move towards a Medical Foundation Model for neuroimages.
In addition, BrainSegFounder’s training scheme and model provide a clear advantage over SAM and MedSAM, two powerful existing segmentation foundation models. (1) While SAM is restricted to 2D RGB images, BrainSegFounder is designed to handle 3D medical images with any number of channels as input, providing greater versatility in medical imaging applications. (2) MedSAM requires bounding-box input prompts and its 3D functionality is limited to manually uploading each image to a plugin for prompting and slice-by-slice annotation. Both methods require manual input. In contrast, our model eliminates the need for such manual interventions once trained, streamlining the segmentation process. (3) Although SAM is capable of automated segmentation without input, it lacks the ability to specify a fixed number of classes and instead generates an arbitrary number of classes; this property leads sub-optimally on medical images with specific segmentation tasks (e.g., lesion detection), and cannot be used without additional human input. (4) Neither SAM nor MedSAM efficiently process multimodal data as they generate predictions from a single scan, whereas BrainSegFounder is designed to integrate multiple scans from the same individual.
@@
However, it’s important to note that while BrainSegFounder shows promise in brain tumor segmentation and brain region segmentation, its application in other neuroimaging tasks remains to be explored. One such task is brain tissue segmentation - a common task in automated analysis. Future research should investigate its adaptability to other neurological conditions, its performance in different clinical environments, and its usefulness in additional common analysis tasks.
@@
In conclusion, BrainSegFounder is a significant step forward 3D foundation models for medical image segmentation and analysis, particularly for multi-modal neuroimaging. Its development underscores the potential of AI and foundation models in enhancing diagnostic accuracy and efficiency, paving the way for more advanced, adaptable, and robust AI tools in healthcare.
@@
This work was partially supported by the National Science Foundation (1908299, 2318984, 2123809) and the National Institute of Aging of the National Institutes of Health (RF1AG071469). This research has been conducted using the UK Biobank Resource under application number 48388. We extend our appreciation to the contributors of the UK Biobank, BraTS, and ALTAS datasets, whose extensive data collection efforts have been fundamental to the success of this research. We gratefully acknowledge NVIDIA AI Technology (NVAITC) ’s support for this research project, especially on GPU parallelization technology. Special thanks to the University of Florida’s HiPerGator-AI supercomputer team for providing the computational resources essential in training the BrainSegFounder models. We appreciate the support from Ying Zhang (UFIT Research Computing) and Kaleb Smith (NVAITC) in providing the necessary computing resources and technology support.
@@ 5.1 UK Biobank Data
Table 8 provides fold-wise comparison of our BrainSegFounder models across all tested parameters.
SwinUNETR (62M) BSF-T (62M) BSF-S (64M) BSF-B (69M) One-Stage (62M) One-Stage (64M) One-Stage (69M)
Fold 1 0.8854 0.9027 0.9032 0.9014 0.9019 0.8994 0.8999
Fold 2 0.9059 0.9181 0.9182 0.9164 0.9188 0.9186 0.9055
Fold 3 0.8981 0.9102 0.9121 0.9097 0.9119 0.9125 0.9002
Fold 4 0.8924 0.9103 0.9100 0.9070 0.9107 0.9133 0.9109
Fold 5 0.9035 0.9139 0.9141 0.9101 0.9132 0.9114 0.9103
Average 0.8971 0.9110 0.9115 0.9089 0.9112 0.9110 0.9054
Table 8: Comparison of BrainSegFounder models through 5-fold cross-validation with metric Dice coefficient on BraTS. SwinUNETR is the winning solution on BraTS challenge 2021, which is performed with fully supervised learning without UKB pretraining. BrainSegFounder is the proposed method, which is conducted with the two-stage pretraining and then finetuning on the target dataset. The one-stage means that pretraining on UKB is performed but not on the BraTS. Note: The performance results for SwinUNETR were published on the official GitHub, utilizing hyper-parameter settings similar to those in our finetuning stage but without implementing the ensembling approach that was described in the published work.
Table 9 presents a comparison of few-shot learning Dice scores on the testing set at varying levels of input training data.
Repeat 5% 10% 20% 40%
BSF SwinU BSF SwinU BSF SwinU BSF SwinU
1 0.7810 0.7437 0.8771 0.8594 0.8893 0.8837 0.8949 0.8885
2 0.7912 0.7899 0.8764 0.8553 0.8814 0.8834 0.8981 0.8981
3 0.7797 0.7886 0.8683 0.8812 0.8901 0.8869 0.8906 0.8956
4 0.8071 0.7966 0.8705 0.8735 0.8911 0.8844 0.9048 0.8938
5 0.7758 0.7893 0.8869 0.8656 0.8895 0.8860 0.8857 0.8938
Average 0.7870 0.7816 0.8758 0.8670 0.8883 0.8849 0.8948 0.8937
Table 9: Comparison of BrainSegFounder (BSF) and SwinUNETR (SwinU) Baseline models trained on 5 repeats of varying percentages of the input data. Data was randomly sampled from the BraTS training dataset, and models were evaluated on the testing dataset.